# Simulating jet launching in binary neutron star mergers with no artificial atmosphere limitations

### v1

@@ III Results
We further note that rest-mass densities within the incipient jet in cases A and C drop below $\!\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3, implying that a uniform density floor with $\rho_{\mathrm{atm}}\!=\!\rho^{*}$ would have affected the evolution. Going to larger time and spatial scales, after the jet breaks out of the surrounding environment and starts to accelerate to high Lorentz factors, the limitations of a high atmospheric floor would become prohibitive. In order to describe jets that are physically consistent and, at the same time, able to evolve into high-$\Gamma$ outflows compatible with GRBs, it is mandatory that their rest-mass densities remain always above the floor. In this work, we have successfully conducted simulations where this condition is strictly maintained at all times.
In case B, the maximum radial velocities never reach values beyond $0.3\,c$ (Figure 4), despite the factor $\sim\!100$ higher magnetic energy and factor $\sim\!3$ higher $L_{\rm EM}$ compared to cases A and C (Figure 3). Here, the fastest part of the outflow remains $2-3$ orders of magnitude more dense (Figure 4) and thus the very low density floor does not make the difference as for the other cases. The final outcome, similar to what found in [11], corresponds to a slow and heavy outflow that has no chance to evolve into a GRB-like jet. Therefore, employing a much lower and more realistic atmospheric floor further favors (in BNS merger context) a GRB progenitor scenario in which the jet is launched by a BH-disk central engine.
@@ IV Conclusions
In this work, we performed BNS merger simulations with the GRMHD code Spritz, considering cases where the remnant massive NS is either long-lived or collapses into a BH $\simeq\!25$ ms after merger. In the latter scenario, an incipient jet emerges, with terminal Lorentz factor and Poynting-flux luminosity that are compatible with a GRB jet. This result, similar to what obtained in previous work (e.g., [7, 8]), is however not sufficient to claim that the system will be able to power a GRB. Further evolution is necessary, knowing that the fate of the incipient jet will strongly depend on the interaction with the surrounding environment.
When exploring the jet propagation on increasingly large scales, densities within the outflow rapidly decrease and the presence of an atmospheric floor (i.e. a minimum density imposed in the simulation) will eventually become a major limitation, unless the floor level is low enough to always remain well below the system’s physical densities. In the jet-forming BNS merger simulations presented here, the latter condition is fulfilled.
While the evolution only covers about $50-75$ ms after collapse, the velocities reached in our simulations ($0.7-0.8\,c$) show the potential advantage of employing a much lower and more realistic density floor, also in comparison to similar models in the literature. In absence of collapse, on the other hand, the very low floor has minor impact on the emerging polar outflows, which are much denser than the incipient jets of BH origin. In this case, velocities remain limited to $\lesssim\!0.3\,c$, further favoring BHs as GRB central engines.
Different caveats affect our simulations, including the low resolution (and consequent unrealistically strong pre-merger magnetic fields) and the use of simple piecewise-polytropic EOS without neutrino transport. Nonetheless, overcoming the limitations associated with the density floor, accomplished here, is independent from the above aspects and represents one necessary step toward a unified description connecting jet-forming BNS mergers with their GRB-related observational signatures.

### v2

@@ III Results
We further note that rest-mass densities within the incipient jet in cases A, C and D drop below $\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3, implying that a uniform density floor with $\rho_{\mathrm{atm}}\!=\!\rho^{*}$ would have affected the evolution. Going to larger time and spatial scales, after the jet breaks out of the surrounding environment and starts to accelerate to high Lorentz factors, the limitations of a high atmospheric floor would become prohibitive. In order to describe jets that are physically consistent and, at the same time, able to evolve into high-$\Gamma$ outflows compatible with GRBs, it is mandatory that their rest-mass densities remain always above the floor. Here, we have successfully conducted simulations where this condition is strictly maintained at all times.
@@ IV Conclusions
In this work, we have carried out simulations of BNS mergers with the GRMHD code Spritz, considering cases where the resulting MNS collapses into a BH about 25 ms or 50 ms after merger, plus one case with no collapse. In the collapsing cases, the system naturally produces two distinct, sequential outflows: a pre-collapse, magnetically driven polar outflow from the MNS and a post-collapse incipient jet powered by the accreting BH. The interaction between these two outflow components, leading to shock heating and potentially responsible for specific EM signatures, is self-consistently captured and discussed for the first time.
All the emerging incipient jets show properties at the base and terminal Lorentz factors compatible with a GRB. Nonetheless, changing the collapse time from $\simeq\!25$ ms to $\simeq\!50$ ms has a major impact on the jet’s ability to pierce through the environment, the latter being significantly more massive and extended in the second case. We conclude that the collapse time is a key parameter to determine the final properties of the escaping jet, as well as the timescale to break out of the dense MNS-driven outflow. Such a result offers a promising coherent framework for interpreting jet diversity in BNS mergers: very different outcomes, from fully choked outflows to successful GRB-like jets, could be explained by the ample range of possible MNS lifetimes before a BH-driven incipient jet is launched.
In our longest simulation (case C), which extends to $\simeq\!135$ ms post-collapse, the BH-powered jet gets close to breaking out of the pre-collapse outflow, with its front reaching spatial scales of $\sim\!10^{4}$ km. The jet advancement through the surroundings is not continuous, but characterized by stalling episodes. This results from the combination of a discontinuous energy injection from the BH-disk engine and the interaction with a non-uniform and evolving environment. Such an intermittent behaviour will likely be imprinted in the final GRB signal, and it may represent a dominant factor governing its variability.
In the no-collapse model, the MNS drives a steady, collimated outflow that remains too dense and too slow to possibly evolve into GRB-like jet. This result strongly favors a BH origin for GRBs produced in BNS mergers.
Figure 5: Rest-mass density profiles along the z-axis (south side) for cases C and D, at 1 ms or 26 ms after BH formation.
A key technical advance in our investigation is the use of an unprecedentedly low density floor (radially declining as $\rho_{\rm atm}\!\propto\!r^{-6}$), which prevents artificial effects on all physically relevant outflows (dynamical ejecta, pre-collapse MNS wind, and BH-powered incipient jet) over the simulated spatial and temporal scales. For the ultimate goal of connecting the jet launching dynamics and early propagation to the final EM signatures and thus the GRB phenomenology, the absence of floor effects is a necessary condition.
Different caveats still affect our simulations, including the limited resolution (and consequent need for unrealistically strong pre-merger magnetic fields) and the use of a simple piecewise-polytropic EOS without neutrino transport. While overcoming such limitations is a central goal of our future work, this will not affect the main conclusions on the two distinct (MNS- and BH-driven) outflows, the role of the collapse time, or the advantage of a BH central engine in powering GRB jets.
@@ Acknowledgements.
V Supplemental Material
V.1 Collapse time and jet launching
The survival time of the MNS remnant determines the duration of the pre-collapse polar outflow and the level of baryon loading in the funnel along the spin axis of the newly-formed BH. Figure 6 shows a meridional view of rest-mass density around collapse time for cases C and D (at $24$ and $51$ ms after merger, respectively). A delay of only $\approx\!25$ ms already produces very different conditions around the merger site. In case C, nascent polar outflows have just started to emerge and extend to $\sim\!500$$700$ km, while in case D the same outflows have expanded to $\sim\!2000$ km and dominate the mass loading along the $z$-axis. In the latter case, while the outflow front is expanding, we see that a magnetized, overpressured cocoon is inflated, resulting in a more difficult propagation environment for the BH-driven incipient jet launched afterwards.
Figure 6: Snapshots of rest-mass density on the xzxz-plane around to the collapse time of the MNS remnant, i.e. at $24$ and $51$ ms after merger for cases C (left) and D (right), respectively.
Figure 7: Meridional view of the rest-mass density, $u_{t}$, internal energy density $\rho\epsilon$ and effective specific entropy $s_{\rm eff}$ for cases C (top panels) and D (bottom panels) at $\approx\!25$ ms after BH formation.
After collapse, the emerging jet encounters the MNS-driven polar outflow and makes an effort to drill through and eventually break out. Figure 7 compares this interaction $\approx\!25$ ms after BH formation for cases C (top) and D (bottom) in terms of rest-mass density $\rho$, $u_{t}$ (a proxy for unbound matter), internal energy density $\rho\epsilon$, and effective specific entropy $s_{\rm eff}\!\equiv\!\ln(P/\rho^{\gamma_{\rm eff}})$, where $\gamma_{\rm eff}\!\equiv\!P/(\rho\epsilon)+1$ (indicative of shock heating).
In both cases, the jet emerges earlier on the south side (z<0z\!<\!0). This asymmetry likely results from a north/south difference in the pre-collapse outflow morphology, common to the two cases C and D. Focusing on the south side, we see how the denser funnel in case D forces the jet to spend more energy and time to propagate, implying significantly different breakout timescale and final properties of the escaping jet.
As already emphasized in the main text, jet’s energy dissipation is accompanied by shock heating, well visible in $\rho\epsilon$ panels. For case C, south side, this effect is particularly evident also looking at $s_{\rm eff}$. These shocks can leave a unique signature in the EM emission, possibly producing a precursor signal to the short GRB or directly contributing to the latter.
Figure 8: Magnetization $b^{2}/(2\rho)$ on xz-plane for cases A-D at 75 ms after merger (first four panels), and for case C at $\approx\!50$ ms after merger (last panel).
We next examine the magnetization near the BH or MNS remnant as a proxy for the acceleration potential of the incipient jet. In particular, Figure 8 shows the magnetic-to-rest-mass energy density ratio $b^{2}/(2\rho)$ for cases A–D at 75 ms after merger and, only for case C, also at about 50 ms after merger. When this quantity is much larger than unity, it corresponds approximately to the terminal Lorentz factor $\Gamma_{\infty}$.
Close to the BH, cases A and C reach values above $10^{4}$. Comparing cases C and D at about $25$ ms after collapse (two rightmost panels in Fig. 8), case C exhibits orders of magnitude higher magnetizations on the south side, consistent with a cleaner funnel and an incipient jet of smaller density (see Figure 7). Finally, we also note that the non-collapsing case B does not show a magnetization capable of producing significant acceleration, suggesting that the polar outflow will remain rather slow also at increasingly large distances.
Overall, the comparison between cases C and D shows the strong impact of the collapse time in the context of jet launching. As concluded in the main text, we find that an earlier collapse limits baryon pollution in the funnel and favors a faster jet propagation which is also energetically less expensive; later collapse allows the pre-collapse MNS-driven outflow to inflate a heavier and more extended cocoon, delaying the jet breakout and dimming the EM emission powered by the jet itself. A denser and more massive cocoon should also enhance photospheric emission and delay high-energy transparency, potentially offering multiwavelength diagnostics of the collapse time.
V.2 Numerical floor prescription
As described in the main text, in all our simulations we employ a radially decreasing numerical rest-mass density floor based on the power-law profile $\rho_{\mathrm{atm}}\!=\!\rho^{*}\,(r/r^{*})^{-6}$ (with $\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3 and $r^{*}\!=\!74$ km). The sixth power of the radial distance, shown in [Pavan2021, Pavan2023] to be a good choice to avoid artificial effects at increasing large scales, corresponds to the lowest numerical floor ever employed in BNS merger simulations. As already noted, radially decreasing profiles where also employed in previous work. For instance, in [hayashi2024jet], the floor scales as $\rho_{\mathrm{atm}}\!\propto\!r^{-3}$ down to 0.166 g/cm3 and then remains constant, giving a floor level higher than ours above $\sim\!700$ km distance and, e.g., $\sim\!1000$ times larger at 2000 km.
Figure 9: Rest-mass density (top) and radial velocity (bottom) along the positive $x$ and $z$ axes at 20 ms after merger, for case A. Rest-mass density floor profiles in the top panel are shown as a black dashed line for our simulations and a light blue dotted line for [hayashi2024jet] (see text).
In Figure 9, top panel, we show the rest-mass density profiles along $x$- and $z$-axis obtained at 20 ms after merger in case A, compared to the floor employed in our work and in [hayashi2024jet]. From such a comparison, we see that the densities of the dynamical ejecta produced during the merging process remain orders of magnitude above our floor, allowing us to correctly describe them up to their natural front with velocities of about 0.9c0.9\,c (Figure 9, bottom panel). Coversely, employing the same floor as in [hayashi2024jet] would have started affecting such a high-velocity tail already at 20 ms after merger, eventually becoming a critical limitation at increasingly large distances.
To further assess the effects of the floor, we also performed an additional simulation (case E) analogous to case C but employing a rest-mass density floor profile as $\rho_{\rm atm}\!\propto\!r^{-2}$. In this case, we only evolved up to 42 ms after merger, i.e. less than 20 ms after the time of collapse to a BH. A meridional view of the rest-mass density comparing cases C and E at such time is shown in Figure 10. At this early evolution stage, we can already observe relevant differences in both the dynamical ejecta and the post-merger outflows. Going to later times, such discrepancies will inevitably grow more and more.
The above results indicate that a floor with $\rho_{\rm atm}\!\propto\!r^{-2}$ is not sufficiently low for a reliable description of the emerging incipient jet and its propagation across the environment.
Figure 10: Meridional view (northern hemisphere only, $z>0$) of the rest-mass density at 42 ms after merger, for case C (left panel) and case E (right panel).
V.3 Step function for $\gamma_{\rm th}$
The simulations presented in this work employ a step function for the adiabatic index $\gamma_{\rm th}$ of the thermal component of the hybrid EOS.
Specifically, for rest-mass densities above $10^{10}\,\text{g/cm}^{3}$ we set $\gamma_{\rm th}\!=\!1.8$, which is a typical value adopted in the literature, while below such densities we impose $\gamma_{\rm th}\!=\!4/3$, corresponding to the ultra-relativistic limit. The latter choice reflects the expectation that, in realistic microphysical EOS, the effective ideal-gas adiabatic index  [^2] This corresponds to the adiabatic index that a simple ideal gas would have for given values of $\rho$, $P$, and $\epsilon$. $\gamma_{\rm eff}\!\equiv\!P/(\rho\epsilon)+1$ becomes 4/3 at low rest-mass densities, as discussed in [Bauswein2010]. The choice of $10^{10}\,\text{g/cm}^{3}$ as threshold value is also consistent with the results of [Bauswein2010] (see Figure 2 therein).
Figure 11: Meridional view of the effective adiabatic index $\gamma_{\mathrm{eff}}$ (top panels; see text for definition) and rest-mass density (bottom panels) at 48 ms after merger for cases C (left) and F (right).
In Figure 11 (top left panel), we present a meridional view of $\gamma_{\rm eff}$ for case C at 48 ms after merger. In the lower left panel, we show instead the rest-mass density. For comparison, we performed an additional simulation (case F) in which we maintained the same adiabatic index $\gamma_{\rm th}\!=\!1.8$ at any density. Case F is also reported in Figure 11 (right panels).
As shown in the upper panels, the value of $\gamma_{\rm th}$ at low densities (i.e. $\rho\!<\!10^{10}\,\text{g/cm}^{3}$) essentially determines the EOS behavior in terms of $\gamma_{\rm eff}$ for both the emerging incipient jet and the environment in front of it. In those regions, $\gamma_{\rm eff}$ is indeed close to 4/3 in case C, and around 1.8 in case F.
Comparing the rest-mass density in the lower panels, we see that the different choice of the adiabatic index at low densities has a strong impact on the system dynamics. In particular, the incipient jet emergence is favored in case F.
In conclusion, (i) introducing the step function for $\gamma_{\rm th}$ allows us to reproduce a realistic adiabatic index of $\approx\!4/3$ for the jet and its propagation environment, while maintaining the same EOS behavior at $\rho\!>\!10^{10}\,\text{g/cm}^{3}$, and (ii) the resulting evolution is rather different from a case where $\gamma_{\rm th}$ is instead constant and equal to 1.8.